### v1

# Algebraic perturbation theory: traversable wormholes and generalized entropy beyond subleading order

Abstract
The crossed product has recently emerged as an important tool in high-energy theory. We combine this with another powerful tool, namely pertubation theory, and study the crossed product algebra of a system under a deformation, relating the structure of deformed observables to that of the undeformed theory. In particular, we derive the change in the von Neumann entropy of the type II algebras, and demonstrate that our approach allows one to formally compute this to arbitrarily high orders in perturbation theory. As a concrete example, we apply this machinery to the case of a double-trace deformation of the thermofield double state in AdS/CFT, which is dual to a traversable wormhole in the bulk, obtaining several new contributions to the generalized entropy relative to the original work by Gao, Jafferis, and Wall Gao:2016bin . We comment on the relevance of this framework for black hole evaporation and interiors, as well as on the applicability of the algebraic approach to quantum gravity more generally.
1 Motivation
Entanglement entropy has played an increasingly important role in holography and black hole physics. The connection between entropy and the area of event horizons, first uncovered by Bekenstein PhysRevD.7.2333 and Hawking PhysRevLett.26.1344 , triggered the longstanding black hole information paradox PhysRevD.14.2460 and contained the germ of the holographic principle Susskind:1994vu ; Bousso:2002ju . Today, one of the most powerful tools in studying the latter is an incarnation of this connection in AdS/CFT, between the entanglement entropy of the boundary field theory and the area of some suitably-defined minimal area surface in the bulk. In analogy to Bekenstein’s relation, this was first proposed by Ryu and Takayanagi in the form Ryu:2006bv
$S=\frac{A_{\gamma}}{4G_{\mathrm{N}}}\leavevmode\nobreak\ ,$ (1)
where $S$ is the entanglement entropy of some subregion $B$ on the boundary, $A_{\gamma}$ is the area of a $d$-dimensional static minimal surface $\gamma$ in AdSd+2 which ends on $\partial B$ (now called the Ryu-Takayanagi or RT surface), and $G_{\mathrm{N}}$ is Newton’s constant. The conjecture (1) was subsequently derived in Lewkowycz:2013nqa using Euclidean methods, and triggered a wave of developments in understanding how to reconstruct bulk physics from the boundary in AdS/CFT (see for example Wall:2012uf ; Dong:2016eik ; Harlow:2016vwg ), thereby also indicating entanglement entropy as an important ingredient in quantum gravity. Among these was the realization that, again in analogy with black hole entropy, subleading corrections from field excitations (e.g., matter, propagating gravitons) must also be included, culminating in the generalized entropy
$S_{\mathrm{gen}}=\frac{A}{4G_{\mathrm{N}}}+S_{\mathrm{matter}}\leavevmode\nobreak\ ,$ (2)
where the area is that of the so-called quantum extremal surface (QES) Engelhardt:2014gca , which may be thought of as the RT surface with the leading quantum corrections included. More recently, these developments have come full-circle as the generalized entropy (2) and the QES have become important ingredients in recent toy models of black hole evaporation in AdS Penington:2019npb ; Almheiri:2020cfm .
However, neither term in (2) is well-defined: in the case of the leading area term, this is due to the need to place a cut-off in the definition of the surface as one approaches the asymptotic boundary, while for the subleading $S_{\mathrm{matter}}$ term, this can be traced to the divergences that arise when attempting to define entanglement entropy in quantum field theory more generally. [^1] See Calabrese:2004eu for a classic review of entanglement entropy in QFT, as well as Rangamani:2016dms ; Headrick:2019eth ; Nishioka:2018khk for reviews which highlight the geometrical relations discussed above, or Hollands:2017dov for a more algebraic treatment. Ultimately, these issues arise from the difficulties of defining entropy in von Neumann algebras of type III, where neither a trace nor reduced density matrices exist. This presents a seemingly fundamental obstacle for QFT, since the algebras of interest are almost always of this type. Phrased more provocatively, the fact that the black hole information paradox exists at all starkly indicates that physicists do not understand entropy as well as we like to think.
In a remarkable recent work Witten:2021unn , Witten proposed a solution to this problem utilizing the crossed product. The basic idea is to consider the type II algebra that results upon adjoining the generator of the modular automorphism group to the original type III algebra Takesaki1973 . Unlike algebras of type III, those of type II do admit a trace, thereby enabling the definition of reduced density matrices and, most interesting for our purposes, a well-defined von Neumann entropy, ${S=-\mathrm{tr}\rho\ln\rho}$ (the relation between the von Neumann and generalized entropies will be discussed later in the text). In light of the fundamental role that entanglement entropy appears to play in black holes and AdS/CFT, it is hoped that a better understanding of this quantity via this more rigorous approach may enable us to make further progress towards quantum gravity. [^2] For example, several works Jefferson:2018ksk ; Jefferson:2019jev ; Raju:2021lwh have pointed out that some of the assumptions underlying the more cavalier approach to entropy generally taken in high-energy theory (e.g., tensor product factorization in defining reduced density matrices) may be problematic, again motivating the need for a different approach.
This crossed product construction has since been generalized to a wide range of systems, including a microcanonical subsector of the TFD Chandrasekaran:2022eqq , de Sitter space Chandrasekaran:2022cip , Rindler space, ball-shaped regions in CFTs, and subregions in AdS/CFT AliAhmad:2023etg , and JT gravity Penington:2023dql ; Kolchmeyer:2023gwa , among others Jensen:2023yxy ; Klinger:2023tgi . Importantly however, as emphasized in AliAhmad:2023etg , this construction relies on fixing the algebra of observables (relative to a given background), meaning that said algebra only contains so-called trivial excitations which do not backreact on the spacetime. However, if we are to understand black hole evaporation, much less quantum gravity more generally, we will certainly need to accommodate changes in the background that shift the defining region of the algebra. As we will show in this work, the crossed product approach naturally handles such deformations of the algebra, enabling us to formally compute the resulting change in observables to arbitrarily high orders in the corresponding perturbation. We regard this as a small but important step towards a perturbative approach to quantum gravity in the algebraic framework. [^3] Of course, there is a rich history within the mathematical physics literature of attempting to synthesize gravity with the operator algebraic approach, however unfortunately aimed primarily at mathematicians; see for example RejznerBook ; Brunetti:2013maa ; Haag:1990ht for complementary works in this vein. In light of the new appreciation for these approaches within high-energy theory, we hope to see greater cross-talk between our two communities in the future.
As an application, we then apply the general perturbative framework we develop to the case of a double-trace deformation of the thermofield double (TFD) state. As shown by Gao, Jafferis, and Wall (GJW) Gao:2016bin , this is dual to a traversable wormhole in the bulk, which causally connects the left and right exteriors of the eternal black hole in AdS for a time set by the strength of the deformation $h$. In that paper, the authors work to linear order in $h$, and argue for a single correction to the subleading $S_{\mathrm{matter}}$ term in the generalized entropy (2). Here, we show how the more rigorous approach from the crossed product subsumes and extends this result. Specifically, we obtain three additional terms that contribute at linear order in $h$, including one at $O(1/N)$ and two at $O(1/N^{2})$, and eight terms that contribute at $O(h^{2})$. Contrary to GJW, we find that two of the terms at linear order in $h$ represent changes to the area term in the generalized entropy (as well as five of the $O(h^{2})$ terms). More abstractly, this demonstrates that the crossed product allows for a controlled treatment of the inherent non-locality of interior operators, in a tractable scenario where everything is in principle computable.
The rest of this paper is organized as follows: in sec. 2, we briefly review how one can define von Neumann entropy via the crossed product construction as presented in Witten:2021unn . Section 3 then introduces our main result, namely a general discussion of how a unitary transformation of the original type III algebra may lead to a change in the type II von Neumann entropy of a subalgebra. At the level of observables, the change takes the form of a weighting of expectation values when computing the trace, which we show can be computed perturbatively to, in principle, arbitrarily high orders in the strength of the perturbation. Then in sec. 4, we apply this general machinery to the example of a double-trace deformation of the traversable wormhole as introduced in GJW Gao:2016bin , and show how our approach allows us to go significantly beyond their results, obtaining eleven new corrections to the generalized entropy up to quadratic order in the strength of the deformation. We close with a discussion of our results in the broader context of quantum gravity and black holes in sec. 5.
2 Crossed products in AdS-Schwarzschild
Here we briefly review the identification of the type II von Neumann factor of observables on either exterior of the eternal AdS-Schwarzschild black hole; we refer the interested reader to Witten:2021unn for details. For conciseness, we are temporarily ignoring issues associated with normalization in the large-$N$ limit, which we postpone to our application in sec. 4 below. [^4] In addition to facilitating a more streamlined review, the main reason for this is that these subtleties are specific to the canonical ensemble in large-$N$ gauge theories, rather than a general feature of the crossed product construction per se AliAhmad:2023etg . As noted above, this construction has since been generalized to many situations, see for example Chandrasekaran:2022eqq ; Chandrasekaran:2022cip ; AliAhmad:2023etg ; Jensen:2023yxy ; Klinger:2023tgi ; Penington:2023dql ; Kolchmeyer:2023gwa , but we have chosen to review it in the context of the eternal black hole to more readily frame the application in later sections.
Let us take $\mathfrak{A}_{r}$ to be the type III1 algebra in the right exterior, encoding local physics of matter and propagating gravitons. The key is to adjoin the generator of modular time evolution $h_{r}$ to this algebra. Formally, this is the one-sided modular Hamiltonian, so that the generator of time translation invariance in the bulk is
$h=h_{r}-h_{l}\leavevmode\nobreak\ .$ (3)
However, the one-sided generators $h_{r,l}$ do not exist as well-defined operators; intuitively, this is because acting with either one alone would generate singular states at the bifurcation surface. For this reason, we will follow AliAhmad:2023etg in referring to the one-sided modular Hamiltonians as modular charges, to avoid confusion with the true modular Hamiltonian (3). Nonetheless, one can effectively adjoin bounded functions of this charge to the algebra $\mathfrak{A}_{r}$ by instead adjoining $T+X$, where $T=h$ is the generator of the canonical outer automorphism group, and $X=h_{l}$ belongs to the commutant. [^5] This still generates singular states in the left algebra, but that is irrelevant for the purposes of the right exterior. Thus, adjoining $T+X$ is formally equivalent to adjoining $h_{r}$. In particular, one constructs the crossed product
$\widehat{\mathfrak{A}}_{r}\coloneqq\mathfrak{A}_{r}\rtimes\mathbb{R}\leavevmode\nobreak\ ,$ (4)
where $\mathbb{R}$ is the type I factor isomorphic to the automorphism generated by $T$. It is then a standard result in the theory of operator algebras that $\widehat{\mathfrak{A}}_{r}$ is type II Takesaki1973 . Physically, $h_{r}$ generates a time shift on the boundary, which is dual to the ADM mass. In this particular case, one can therefore think of $\widehat{\mathfrak{A}}_{r}$ as the type II algebra obtained by enlarging the original type III1 algebra $\mathfrak{A}_{r}$ to incorporate large gauge transformations, i.e., gravity. As emphasized in AliAhmad:2023etg however, gravity itself is not intrinsically responsible for the type reduction from III to II; rather, the key is to adjoin the generator of the canonical outer automorphism group, i.e., to consider the algebra together with its intrinsic modular dynamics. Indeed, a similar construction holds on the boundary, where the algebra of exterior operators in the right CFT $\mathfrak{A}_{R}$ is dual to $\mathfrak{A}_{r}$: one can formally adjoin the modular charge $H_{R}$ (dual to $h_{r}$) in the same manner, but this does not admit a gravitational interpretation.
In the present context, the significance of this construction is that, since the algebra $\widehat{\mathfrak{A}}_{r}$ is of type II, one can define a trace and therefore a rigorous definition of von Neumann entropy as follows Witten:2021unn : the crossed product algebra $\widehat{\mathfrak{A}}$ acts on the Hilbert space
^$\widehat{\mathcal{H}}\coloneqq\mathcal{H}\otimes L^{2}(\mathbb{R})\leavevmode\nobreak\ ,$ (5)
where $L^{2}(\mathbb{R})$ is the type I Hilbert space acted upon by the modular charge $X$ (n.b., in what follows, it will be important to remember that $X$ does not act on $\mathcal{H}$). We consider separable states of the form
$\widehat{\Psi}\coloneqq\Psi\otimes g(X)^{1/2}\leavevmode\nobreak\ ,$ (6)
with $\Psi$ the vacuum state in $\mathcal{H}$ and $g(X)$ a state in $L^{2}(\mathbb{R})$. For technical reasons, we assume that $g$ is analytic (cf. (33) below). Then for some (but not necessarily all) $\widehat{a}\in\widehat{\mathfrak{A}}$, we may define the trace
$\mathrm{tr}\widehat{a}=\langle{\widehat{\Psi}}|\widehat{a}K^{-1}|{\widehat{\Psi}}\rangle=\int_{-\infty}^{\infty}\!\mathrm{d}X\,e^{X}\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right>\leavevmode\nobreak\ .$ (7)
where $K$ is the density state for $\widehat{\mathfrak{A}}$, which Witten:2021unn constructs by decomposing the corresponding modular operator as
$\widehat{\Delta}_{\widehat{\Psi}}=\widetilde{K}K\leavevmode\nobreak\ ,$ (8)
where
$K=\Delta_{\Psi}e^{-X}g(T+X)=e^{-(T+X)}g(T+X)\qquad\mathrm{and}\qquad\widetilde{K}=e^{X}g(X)^{-1}\leavevmode\nobreak\ ,$ (9)
are the density states on the algebra $\widehat{\mathfrak{A}}$ and its commutant, respectively, and $\Delta_{\Psi}$ is the modular operator on the original type III algebra $\mathfrak{A}$. One can then define the von Neumann entropy on the type II algebra in the usual way, namely
$S(K)=-\mathrm{tr}\,K\ln K\leavevmode\nobreak\ .$ (10)
More generally, one may consider an excited state $\Phi$ obtained by adding matter to the state $\Psi$; the corresponding density state upon performing the crossed product is then Chandrasekaran:2022eqq ,
$K^{\prime}=\epsilon\,\bar{g}\left(\epsilon(T+X)\right)^{1/2}e^{-\beta X}\Delta_{\Phi|\Psi}g\left(\epsilon(T+X)\right)^{1/2}\leavevmode\nobreak\ ,$ (11)
where $\Delta_{\Phi|\Psi}$ is the relative modular operator and $\epsilon$ is a parametrically small constant which controls the semiclassical behavior of the state. Note that when $\Phi=\Psi$, this reduces to ${\Delta_{\Phi|\Psi}=\Delta_{\Psi}}$, and we recover (9). However, a crucial point is that one cannot remain within the same algebra $\widehat{\mathfrak{A}}$ and consider arbitrarily excited states. The basic reason is that fixing an algebra amounts to fixing a region, but an arbitrary excitation may backreact – shifting the region – or lack a geometric interpretation. As explained in AliAhmad:2023etg , the above holds only for so-called trivial excitations which do not affect the geometry of the region. For the more interesting case in which the QES changes under the perturbation, the expectation value becomes weighted relative to the vacuum state; this is the subject of the present work.
3 Weighting of expectation values under a general perturbation
In this section, we compute the change in the entropy for a type II factor under a unitary perturbation of the original type III1 algebra. In particular, we compute the change in expectation values in the type II crossed product algebra, which become weighted by an exponential factor that depends on the perturbation. As we shall eventually apply this to a double-trace deformation of the thermofield double (TFD) state Gao:2016bin in sec. 4 below, we will have in mind that the unperturbed state $\Psi_{0}$ is the original TFD, while the perturbed state $\Psi$ is the TFD with the double-trace deformation turned on, but the formalism in this section is completely general, and applies to any two type III1 algebras related by some unitary transformation.
Before proceeding to the main computation in subsec. 3.4, we first clarify a potential confusion regarding how entropy can change under a unitary transformation in subsec. 3.1, and verify the corresponding change in the modular Hamiltonian in subsec. 3.2.
3.1 Semifinite partial traces and (non)unitary dynamics
Here, let us explicate what we mean by a change in entropy under a unitary transformation by returning to the more familiar type I algebra of quantum mechanics, where it is a basic fact such a transformation leaves the von Neumann entropy invariant. That is, if we have some density state $\rho$ that transforms as $\rho\mapsto U\rho U^{\dagger}$ under some unitary operator $U$, then the von Neumann entropy $S=-\mathrm{tr}\rho\ln\rho$ remains unchanged:
$S\mapsto-\mathrm{tr}\left[U\rho U^{\dagger}\ln(U\rho U^{\dagger})\right]=-\mathrm{tr}\left[U\rho U^{\dagger}U(\ln\rho)U^{\dagger}\right]=-\mathrm{tr}\rho\ln\rho\leavevmode\nobreak\ ,$ (12)
where the last step relied crucially on the cyclicity of the trace,
$\mathrm{tr}(AB)=\mathrm{tr}(BA)\leavevmode\nobreak\ .$ (13)
However, this implicitly relied on the operators $A,B$ acting on the same Hilbert space on which the trace is defined. To see why this is important, suppose we now consider a factorizable Hilbert space,
$\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}\leavevmode\nobreak\ .$ (14)
The type I algebras acting on $\mathcal{H}_{A}$ and $\mathcal{H}_{\bar{A}}$ both admit a unique partial trace, respectively $\mathrm{tr}_{A}$ and $\mathrm{tr}_{\bar{A}}$, which allows us to define operators acting on either Hilbert space. The key fact is that these partial traces are only cyclic on their corresponding algebra. As a trivial example, consider an operator $\mathcal{O}=A_{1}A_{2}\otimes\bar{A}_{1}\bar{A}_{2}$, where $A_{i}$ acts on $\mathcal{H}_{A}$, and $\bar{A}_{i}$ acts on $\mathcal{H}_{\bar{A}}$. We may define the reduced operator acting only on $\mathcal{H}_{A}$ via the trace on the complement,
$\mathrm{tr}_{\bar{A}}\mathcal{O}=A_{1}A_{2}\,\mathrm{tr}\left(\bar{A}_{1}\bar{A}_{2}\right)=A_{1}A_{2}\,\mathrm{tr}\left(\bar{A}_{2}\bar{A}_{1}\right)\leavevmode\nobreak\ ,$ (15)
since the trace over the complement does not act on $\mathcal{H}_{A}$. [^6] If $[A_{1},A_{2}]=0$, then one obtains the same result as if the partial trace were cyclic on the full Hilbert space, but the latter is nonsensical since it does not even act on both factors. More pertinently, this means that if we consider a unitary transformation defined on the full Hilbert space, then
$\mathrm{tr}_{\bar{A}}\left(U\mathcal{O}U^{\dagger}\right)\neq\mathrm{tr}_{\bar{A}}\left(U^{\dagger}U\mathcal{O}\right)=\mathrm{tr}_{\bar{A}}\,\mathcal{O}\leavevmode\nobreak\ ,$ (16)
since we cannot apply the cyclic property of the trace over $\mathcal{H}_{\bar{A}}$ to an operator $U$ with support on $\mathcal{H}_{A}$. Only in the special case where $U$ is defined only on $\mathcal{H}_{\bar{A}}$ would (16) hold.
In other words, unitary evolution on $\mathcal{H}$ generally does not descend to unitary evolution on $\mathcal{H}_{A}$. One can of course still define a density state on the subsystem after unitary evolution on the whole,
$\rho^{\prime}_{A}=\mathrm{tr}_{\bar{A}}\left[U(\rho_{A}\otimes\rho_{\bar{A}})U^{\dagger}\right]\neq U\rho_{A}U^{\dagger}\leavevmode\nobreak\ ,$ (17)
but it does not make sense to invoke cyclicity here, since this partial trace is defined on $\mathcal{H}_{\bar{A}}$ only, whereas $U$ has support on the full Hilbert space. Rather, the process of going from $\rho_{A}$ to $\rho^{\prime}_{A}$ is generally described by a quantum channel instead of a unitary operator,
$\rho^{\prime}_{A}=\sum_{i}E_{i}\rho_{A}E_{i}^{\dagger},$ (18)
where $E_{i}$ are Kraus operators. If there is one element in the sum, then this reduces to unitary evolution on the subsystem, but it will not be the same as $U\rho_{A}U^{\dagger}$ for the global unitary $U$ defined above.
Returning to our case of interest, it is impossible to define a partial trace on either the algebra or its commutant because both are type III1. For that reason, we invoke the crossed product construction to obtain a semifinite type II factor, which allows us to define a trace. As in the type I case just reviewed however, this trace will not be cyclic for a unitary operator acting on both the algebra and its commutant, so the global unitary transformation will still lead to non-trivial changes in density operators and expectation values defined on the subsystem. This is what allows the von Neumann entropy of the type II algebra (e.g., of the right exterior in the TFD) to change under the interaction Hamiltonian.
3.2 Covariant modular structure
Here, we establish some notation that will be used throughout the remainder of the paper, and in particular verify that the modular Hamiltonian transforms as expected under the perturbation. Let $\mathfrak{A}_{0}$ be a type III1 algebra acting on $\mathcal{H}_{0}$, and $\widehat{\mathfrak{A}}_{0}$, ^$\widehat{\mathcal{H}}_{0}$ the corresponding objects obtained by adjoining the modular charge $X_{0}$ as per the crossed product construction above. Similarly, let $\mathfrak{A}$ by the type III1 algebra algebra obtained from $\mathfrak{A}_{0}$ by applying a unitary transformation,
$\mathfrak{A}=U\mathfrak{A}_{0}U^{\dagger}\leavevmode\nobreak\ ,$ (19)
which acts on $\mathcal{H}$; the corresponding crossed product objects will be denoted $\widehat{\mathfrak{A}}$, $\widehat{\mathcal{H}}$. Since the unitary does not act on the adjoined $L^{2}(\mathbb{R})$ factor in the crossed product, it follows immediately that $\widehat{\mathfrak{A}}=U\widehat{\mathfrak{A}}_{0}U^{\dagger}$. Note that when applied to the TFD below, $\mathfrak{A}_{0}$ and $\mathfrak{A}$ will denote the full exterior algebra on both sides, whose corresponding type II algebras admit a decomposition of the form $\widehat{\mathfrak{A}}_{0}=\widehat{\mathfrak{A}}_{L,0}\otimes\widehat{\mathfrak{A}}_{R,0}$, and similarly for $\widehat{\mathfrak{A}}$, where the subscripts ,${}_{L},{}_{R}$ denote the left and right sides of the TFD (i.e., $\widehat{\mathfrak{A}}_{L,0}^{\prime}=\widehat{\mathfrak{A}}_{R,0}$).
Associated to $\mathfrak{A}_{0}$ is a cyclic and separating state $|{\Psi_{0}}\rangle$, excitations about which generate the Hilbert space $\mathcal{H}_{0}$, and a modular Hamiltonian $H_{\mathrm{mod},0}$ that annihilates this ground state, $H_{\mathrm{mod},0}|{\Psi_{0}}\rangle=0$. Acting with the unitary (19) implies that the corresponding cyclic and separating state for $\mathfrak{A}$ is $|{\Psi}\rangle=U|{\Psi_{0}}\rangle$. That $|{\Psi}\rangle$ has these properties follows immediately from those of the original state, since
$\mathfrak{A}|{\Psi}\rangle=(U\mathfrak{A}_{0}U^{\dagger})U|{\Psi_{0}}\rangle=U\mathfrak{A}_{0}|{\Psi_{0}}\rangle\leavevmode\nobreak\ .$ (20)
Now, if we have access to density operators – which we do for the type II algebras $\widehat{\mathfrak{A}}_{0}$ and $\widehat{\mathfrak{A}}$ – then this suggests that $\widehat{\rho}_{0}=|{\widehat{\Psi}_{0}}\rangle\langle{\widehat{\Psi}_{0}}|$ also transforms to $\widehat{\rho}=U\widehat{\rho}_{0}U^{\dagger}$. [^7] Note that this does not contradict (17) because here, $\rho$ is the density operator on the full system, not one obtained via a partial trace. Formally, if we think of $\rho=e^{-H_{\mathrm{mod}}}$, then this suggests that the modular operator transforms as
$H_{\mathrm{mod}}=UH_{\mathrm{mod},0}U^{\dagger}\leavevmode\nobreak\ .$ (21)
That this preserves the separating property of the vacuum state again follows immediately from the unperturbed state:
$H_{\mathrm{mod}}|{\Psi}\rangle=(UH_{\mathrm{mod},0}U^{\dagger})U|{\Psi_{0}}\rangle=UH_{\mathrm{mod},0}|{\Psi_{0}}\rangle=0\leavevmode\nobreak\ .$ (22)
A more significant requirement is that it generates the corresponding modular automorphism group, i.e., that $|{\Psi}\rangle$ is a KMS state relative to $H_{\mathrm{mod}}$. Defining $M(s)\coloneqq\Delta^{-is}$ where ${\Delta\coloneqq e^{-H_{\mathrm{mod}}}}$ is the modular operator for the deformed TFD, the KMS condition reads Haag:1996hvx
${\Psi}\right|{M(t)aM(-t)b}\left|{\Psi}\right>=\left<{\Psi}\right|{bM(t+i\beta)aM(-t-i\beta)}\left|{\Psi}\right>\leavevmode\nobreak\ ,\qquad\forall a,b\in\mathfrak{A}\leavevmode\nobreak\ .$ (23)
Equivalently, since $H_{\mathrm{mod}}$ annihilates the state,
${\Psi}\right|{aM(-t)b}\left|{\Psi}\right>=\left<{\Psi}\right|{bM(t+i\beta)a}\left|{\Psi}\right>\leavevmode\nobreak\ .$ (24)
Using the fact that $|{\Psi}\rangle=U|{\Psi_{0}}\rangle$ and $M=UM_{0}U^{\dagger}$ (with $M_{0}(s)$ defined in obvious analogy to $M(s)$), this becomes
${\Psi_{0}}\right|{U^{\dagger}aUM_{0}(-t)U^{\dagger}bU}\left|{\Psi_{0}}\right>$ =${\Psi_{0}}\right|{U^{\dagger}bUM_{0}(t+i\beta)U^{\dagger}aU}\left|{\Psi_{0}}\right>$ (25)
${\Psi_{0}}\right|{a_{0}M_{0}(-t)b_{0}}\left|{\Psi_{0}}\right>$ =${\Psi_{0}}\right|{b_{0}M_{0}(t+i\beta)a_{0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ ,$
where we used the fact that, since the algebras are related by $\mathfrak{A}=U\mathfrak{A}_{0}U^{\dagger}$, $a_{0}\in\mathfrak{A}_{0}$ maps to $a\coloneqq Ua_{0}U^{\dagger}\in\mathfrak{A}$, and hence $U^{\dagger}aU=a_{0}\in\mathfrak{A}$; similarly for $b$. Thus, the KMS condition in the new algebra is equivalent to the KMS condition in the original algebra.
We note in passing that we may also define a modular conjugation $J$ for the new algebras in terms of the original one, $J_{0}$, by acting with this same unitary as $J=UJ_{0}U^{\dagger}$. Starting from an element $a_{R}\in\mathfrak{A}_{R}$, we then have
$Ja_{R}J=UJ_{0}U^{\dagger}a_{R}UJ_{0}U^{\dagger}=UJ_{0}\,a_{R,0}\,J_{0}U^{\dagger}=Ua_{L,0}U^{\dagger}=a_{L}\in\mathfrak{A}_{L}\leavevmode\nobreak\ .$ (26)
In short, the modular structure of the original algebra $\mathfrak{A}_{0}$ carries over as expected to the deformed algebra $\mathfrak{A}$.
3.3 General change in weighting of expectation values
Preliminaries thus dispensed with, we can now proceed to our main interest: the change in the weighting of expectation values in the type II theory under a unitary perturbation. As reviewed above, the type III systems $\mathfrak{A}_{0}$, $\mathfrak{A}$ do not admit a tensor product factorization, but the type II systems $\widehat{\mathfrak{A}}_{0}$, $\widehat{\mathfrak{A}}$ do. Accordingly, suppose we factorize these algebras into left ($\widehat{\rho}_{L,0}$, $\widehat{\rho}_{L}$) and right ($\widehat{\rho}_{R,0}$, $\widehat{\rho}_{R}$) density matrices. In the application to the traversable wormhole below for example, $\widehat{\rho}_{R,0}$ and $\widehat{\rho}_{R}$ will be the density matrices of the right exterior (that is, outside the black hole) algebra before and after the deformation, respectively. While the transformation (19) is unitary on the full algebra, it does not necessarily act unitarily on either of these factors as explained above, which allows for the possibility of a change in the von Neumann entropy. Since this can only be rigorously defined in the type II theory, we wish to compute
$\Delta S=S(\widehat{\rho}_{R})-S(\widehat{\rho}_{R,0})\leavevmode\nobreak\ ,$ (27)
with the von Neumann entropies on the right-hand side given by (10), with $K\in\{\widehat{\rho}_{R},\widehat{\rho}_{R,0}\}$.
However, doing this requires that we make more precise some notational liberties taken in the previous crossed product works reviewed above. In particular, note that in (7), the expectation value ${\Psi}\right|{\widehat{a}}\left|{\Psi}\right>$ is an operator-valued function of the modular charge $X$. But as an operator, it is not meaningful to integrate this over the reals as written. Intuitively, what is meant is that we integrate over the (continuous) spectrum of $X$, whose domain is $\mathbb{R}$. This is relevant here because, as we will see, the weighting from the perturbation results in a factor containing commutators of operators which is non-diagonal in the energy eigenbasis, so it is helpful to make this slightly cavalier step to c-numbers more explicit. To that end, observe that when we write
$\widehat{\Psi}=\Psi\otimes g(X)^{1/2}$ (28)
(which is (6), reproduced here for convenience), the second factor is the wavefunction on $L^{2}(\mathbb{R})$. Working in the energy eigenbasis, this is [^8] Any basis will do, but we will eventually want to identify $X$ with the modular charge $H_{L}$, so the energy eigenbasis is the most natural choice.
$g^{1/2}(\lambda)=\langle\lambda|g^{1/2}\rangle\leavevmode\nobreak\ ,$ (29)
where $X|{\lambda}\rangle=\lambda|{\lambda}\rangle$, and $|{g^{1/2}}\rangle$ is the state of the $L^{2}(\mathbb{R})$ factor. The statement that $g^{1/2}(X)$ is an element of $L^{2}(\mathbb{R})$ then becomes
$\int\!\mathrm{d}X\,g(X)\coloneqq\int\!\mathrm{d}\lambda\,|\langle\lambda|g^{1/2}\rangle|^{2}=1\leavevmode\nobreak\ ,$ (30)
where we define the non-rigorous expression on the far left-hand side by the expression in the middle. The physical interpretation of this is that $|\langle\lambda|g^{1/2}\rangle|^{2}$ is the probability of the energy eigenstate $\lambda$ given the prepared state $g^{1/2}$. A correlation function in $L^{2}(\mathbb{R})$ is similarly computed by inserting a complete basis of these eigenstates. In particular, given an arbitrary function $f(X)$,
$f(X)\right>_{L^{2}(\mathbb{R})}=\langle{g^{1/2}}|f(X)|{g^{1/2}}\rangle=\int\!\mathrm{d}\lambda\langle{g^{1/2}}|f(X)|{\lambda}\rangle\langle\lambda|g^{1/2}\rangle=\int\!\mathrm{d}\lambda\,f(\lambda)|\langle\lambda|g^{1/2}\rangle|^{2}\leavevmode\nobreak\ ,$ (31)
since $f(X)|{\lambda}\rangle=f(\lambda)|{\lambda}\rangle$.
Now, in the deformed theory, the state in the crossed product is $|{\widehat{\Psi}}\rangle=|{\Psi}\rangle|{g^{1/2}}\rangle$, so the trace (7) is
$\mathrm{tr}\,\widehat{a}$ =${\widehat{\Psi}}\right|{\widehat{a}K^{-1}}\left|{\widehat{\Psi}}\right>$ (32)
$=\langle{g^{1/2}}|\langle{\Psi}|\widehat{a}\,e^{T+X}g^{-1}(T+X)|{\Psi}\rangle|{g^{1/2}}\rangle$
$=\langle{g^{1/2}}|\langle{\Psi}|\widehat{a}\,e^{X}g^{-1}(X)|{\Psi}\rangle|{g^{1/2}}\rangle$
$=\int\!\mathrm{d}\lambda\langle{g^{1/2}}|\langle{\Psi}|\widehat{a}\,e^{X}g^{-1}(X)|{\lambda}\rangle\langle\lambda|{\Psi}\rangle|{g^{1/2}}\rangle$
$=\int\!\mathrm{d}\lambda\,e^{\lambda}g^{-1}(\lambda)|\langle g^{1/2}|\lambda\rangle|^{2}\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right>$
$=\int\!\mathrm{d}\lambda\,e^{\lambda}\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right>\leavevmode\nobreak\ ,$
as expected. In the third equality, we used the fact that the modular Hamiltonian annihilates the state, $T|{\Psi}\rangle=0$; in the fourth equality, we inserted a complete (energy eigen)basis; the last equality follows from (29). Note that we also used the fact, mentioned below (5), that $X$ acts only on the $L^{2}(\mathbb{R})$ factor.
Relating this to the undeformed expectation value is then quite straightforward: we have
$|{\Psi}\rangle=U|{\Psi_{0}}\rangle\leavevmode\nobreak\ ,\qquad\widehat{a}=U\widehat{a}_{0}U^{\dagger}\leavevmode\nobreak\ ,\qquad K^{-1}=UK^{-1}_{0}U^{\dagger}\leavevmode\nobreak\ ,$ (33)
where the last follows from the assumption below (6) that $g^{-1}$ is analytic. Crucially, note again that since $U$ does not act on $L^{2}(\mathbb{R})$, it commutes with any function of $X$. Hence, inserting the relations (33) in the third line in (32), all factors of $U$, $U^{\dagger}$ commute through and combine to $\mathbb{1}$, so that the last line becomes
[…]
where ${\Psi_{0}}\right|{\widehat{a}_{0}}\left|{\Psi_{0}}\right>$ as well as $\lambda(\lambda_{0})$ is a function of $\lambda_{0}\neq\lambda$. Since the spectral measure is unique, [^9] A unitary transformation preserves eigenvalues and thus must preserve the spectral measure. the Jacobian $\mathrm{d}\lambda/\mathrm{d}\lambda_{0}=1$, so this becomes simply
$\mathrm{tr}\,\widehat{a}=\int\!\mathrm{d}\lambda_{0}\,e^{\lambda(\lambda_{0})}\left<{\Psi_{0}}\right|{\widehat{a}_{0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ .$ (35)
In other words, the expectation value of observables in the perturbed type II algebra is the same as that in the unperturbed algebra up to a weighting factor $e^{\lambda(\lambda_{0})}$. Intuitively, this is the expectation value of $Ue^{X_{0}}U^{\dagger}$ in the energy eigenbasis, where $X_{0}$ denotes the modular charge in the unperturbed theory. To see this, suppose we work directly with the non-rigorous notation in (7), and apply the unitary transformation in (33) to the modular charge (i.e., the one-sided Hamiltonian), i.e., $X=UX_{0}U^{\dagger}$: [^10] Strictly speaking, since $X$, $X_{0}$ act on the $L^{2}(\mathbb{R})$ factor and not on the TFD Hilbert space, we should denote this by some other operator $\widetilde{U}$. This technicality will become clear shortly.
$\mathrm{tr}\,\widehat{a}$ $=\int\!\mathrm{d}X\,e^{X}\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right>$ (36)
$=\int\!\mathrm{d}(UX_{0}U^{\dagger})\,e^{UX_{0}U^{\dagger}}\langle{\Psi_{0}}|U^{\dagger}(U\widehat{a}_{0}U^{\dagger})U|{\Psi_{0}}\rangle$
$=\int\!\mathrm{d}X_{0}\,\left(Ue^{X_{0}}U^{\dagger}\right)\langle{\Psi_{0}}|\widehat{a}_{0}|{\Psi_{0}}\rangle\leavevmode\nobreak\ ,$
where in the last step, we used the fact that a unitary transformation does not alter the spectral measure, and the freedom to move unitary operators in/out of exponentials at will. Since both $U$ and $X_{0}$ are operators however, these must be evaluated in some state [^11] As noted in Witten:2021unn , it does not matter which state we choose, since the factor $g$ will drop out of the final expression. This reflects the fact that the spectrum ultimately depends on the operator, not the state. before the integral over the spectrum can be performed, cf. (30).
Let us make the connection to the weight $e^{\lambda(\lambda_{0})}$ in (34) more explicit. In our application below, we will be interested in the case in which $X$ is identified with the suitably-normalized left modular charge $\widetilde{H}_{L}$, and is related to the unperturbed charge by [^12] The tilde notation concerns the $1/N$ subtleties postponed above, and distinguishes the normalized ${\widetilde{H}_{L}=H_{L}/N}$ vs. unnormalized modular charge $H_{L}$; this will be discussed in subsec. 4.2, and the distinction will be relevant momentarily.
$\widetilde{H}_{L}=\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}\leavevmode\nobreak\ ,$ (37)
where $\delta\widetilde{H}_{L,0}$ contains a parameter that controls the strength of the perturbation, and hence acts as our expansion parameter in this section; we will give a concrete example of this in the application to the TFD in sec. 4. Note however that (37) is exact to all orders, i.e., we are considering
$Ue^{\widetilde{H}_{L,0}}U^{\dagger}=e^{\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}}\leavevmode\nobreak\ ,$ (38)
so that generically $\delta\widetilde{H}_{L,0}$ will contain an infinite series of nested commutators, cf. (53). Now, taking the relevant expectation value from the fourth line of (32), we have
$\langle{\Psi}|\widehat{a}\,e^{\widetilde{H}_{L}}g^{-1}(\widetilde{H}_{L})|{\Psi}\rangle=\langle{\Psi}|\widehat{a}\,\exp\{\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}\}\,g^{-1}(\widetilde{H}_{L}))|{\Psi}\rangle$ (39)
$\approx\langle{\Psi}|\widehat{a}\underbrace{\left[1+\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}+\frac{1}{2}\left(\widetilde{H}_{L,0}^{2}+\delta\widetilde{H}_{L,0}^{2}+\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}\right)\right]}_{[\ldots]}g^{-1}(\widetilde{H}_{L}))|{\Psi}\rangle\leavevmode\nobreak\ ,$
where in the second line we have expanded to second order, and $\{\mathcal{O}_{1},\mathcal{O}_{2}\}$ is the anticommutator of operators $\mathcal{O}_{1}$, $\mathcal{O}_{2}$. In the underbrace, the second-order expansion in square brackets has been denoted []delimited-[][$\ldots$][ … ] for compactness in the sequel. Note that there is no need to expand $g^{-1}(\widetilde{H}_{L})$, because the state of the $L^{2}(\mathbb{R})$ factor above, $|{g^{1/2}}\rangle$, is the state corresponding to the perturbed TFD, so these will eventually cancel as in (32), cf. footnote 11.
To proceed, it is important to note the distinction between $H_{L,0}$ and $\widetilde{H}_{L,0}$: the former is the non-normalized modular charge, which formally acts on the (left factor of the) Hilbert space of the QFT. The latter is a suitably normalized version identified as the charge ${X_{0}=\widetilde{H}_{L,0}}$, which acts only on $L^{2}(\mathbb{R})$. Therefore, the tilded operators in (39) can be moved outside the TFD expectation value. Relatedly, note that the unitary $U$ that deforms the TFD acts only on the QFT Hilbert space, i.e.,
$|{\widetilde{\Psi}}\rangle=U|{\Psi_{0}}\rangle|{g^{1/2}}\rangle\leavevmode\nobreak\ ,$ (40)
so that the $L^{2}(\mathbb{R})$ factor is unaffected. However, the energy eigenstates (denoted $\lambda$ above) are affected, since $H_{L,0}$ – and therefore $X=\widetilde{H}_{L,0}$ – encodes information about the ADM mass, and hence knows about the deformation.
Now, inserting the expectation value (39) back into (32), the trace may be written
$\mathrm{tr}\,\widehat{a}$ $=\int\!\mathrm{d}\lambda\langle{g^{1/2}}|\,\langle{\Psi}|\,\widehat{a}\,[\ldots]|{\Psi}\rangle\,g^{-1}(\widetilde{H}_{L})|{\lambda}\rangle\langle\lambda|g^{1/2}\rangle$ (41)
$=\int\!\mathrm{d}\lambda\langle{g^{1/2}}|[\ldots]\,g^{-1}(\widetilde{H}_{L})|{\lambda}\rangle\langle\lambda|g^{1/2}\rangle\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right>$
$=\int\!\mathrm{d}\lambda\mathrm{d}\lambda^{\prime}\langle g^{1/2}|\lambda^{\prime}\rangle\langle{\lambda^{\prime}}|[\ldots]\,g^{-1}(\widetilde{H}_{L})|{\lambda}\rangle\langle\lambda|g^{1/2}\rangle\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right>$
[…]
where on the third line, we inserted another complete basis, whereupon we exploited the fact that $\langle{\lambda^{\prime}}|f(X)|{\lambda}\rangle=\delta(\lambda^{\prime}-\lambda)f(\lambda)$ since these are the eigenstates of $X$. The final weight ${\lambda}\right|{[\ldots]}\left|{\lambda}\right>\approx\left<{\lambda}\right|{Ue^{X_{0}}U^{\dagger}}\left|{\lambda}\right>$ is precisely the factor of $e^{\lambda}$ in (34), i.e.,
$e^{\lambda(\lambda_{0})}\approx\left<{\lambda}\right|{1+\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}+\frac{1}{2}\left(\widetilde{H}_{L,0}^{2}+\delta\widetilde{H}_{L,0}^{2}+\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}\right)}\left|{\lambda}\right>\leavevmode\nobreak\ .$ (42)
This does not depend on the state of the TFD, which is to be expected since the spectrum of an operator is state-independent. Note that in the final line of (41), $\lambda$ is a dummy variable that we can freely replace with $\lambda_{0}$, reflecting the aforementioned fact that the spectral measure is unique.
However, the expansion (42) is not yet complete: until this point, we have implicitly absorbed the intrinsic temperature associated to the modular group into the modular charge $\widetilde{H}_{L}$, i.e., we have set $\beta=1$ (cf. (23); see for example Haag:1996hvx ; Connes:1994hv for background discussion on this point). Since $\mathfrak{A}\neq\mathfrak{A}_{0}$, we must allow for the possibility that $\beta\neq\beta_{0}$; physically, this is because the QES that acts as the bifurcation point for the boost operator (or the generator of modular flow more generally) may change, and the temperature is set by the associated horizon area. Therefore, we first restore the factors of $\beta$ in (42) arising from the fact that really, $e^{X}=e^{\beta\widetilde{H}_{L}}$:
$e^{\lambda(\lambda_{0})}\approx\left<{\lambda}\right|{1+\beta\widetilde{H}_{L,0}+\beta\delta\widetilde{H}_{L,0}+\frac{1}{2}\beta^{2}\left(\widetilde{H}_{L,0}^{2}+\delta\widetilde{H}_{L,0}^{2}+\{\widetilde{H}_{L,0},\delta\widetilde{H}_{L,0}\}\right)}\left|{\lambda}\right>\leavevmode\nobreak\ .$ (43)
We then take
$\beta=\beta_{0}+\delta\beta_{0}\leavevmode\nobreak\ ,$ (44)
where $\delta\beta_{0}$ includes changes to all orders in the perturbation. In general, $\delta\beta_{0}\sim\delta\widetilde{H}_{L,0}$ in that both of them will be a series of perturbative corrections that, at least in the case of the double-trace deformation below, will be controlled by the same expansion parameter $h$. That is, we may write
$\delta\beta_{0}=\beta_{1}+\beta_{2}+\ldots$ (45)
where $\beta_{i}\sim O(h^{i})$. This implies that in general,
$\delta\beta_{0}\sim\delta\widetilde{H}_{L,0}\sim O(h)\leavevmode\nobreak\ ,$ (46)
to leading order, i.e., each of these in principle contains terms higher-order in $h$. Thus, expanding (43) to second order in $h$, we have
$e^{\lambda(\lambda_{0})}\approx\langle{\lambda}|$ $1+\beta_{0}\widetilde{H}_{L,0}+\frac{1}{2}\beta_{0}^{2}\widetilde{H}_{L,0}^{2}+\,\ldots$ (47)
$+\left(\beta_{0}\,\delta\widetilde{H}_{L,0}+\delta\beta_{0}\widetilde{H}_{L,0}\right)+\frac{1}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\delta\beta_{0}\,\beta_{0}\widetilde{H}_{L,0}^{2}$
$+\frac{1}{2}\beta_{0}^{2}\,\delta\widetilde{H}_{L,0}^{2}+\delta\beta_{0}\left(\delta\widetilde{H}_{L,0}+\beta_{0}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\frac{1}{2}\delta\beta_{0}\widetilde{H}_{L,0}^{2}\right)|{\lambda}\rangle\leavevmode\nobreak\ ,$
where the terms on the first, second, and third line are $O(1)$, $O(h)$, and $O(h^{2})$, respectively. In particular, the ellipsis on the first line indicates the exponential series at $O(1)$, which we can re-sum to obtain
$e^{\lambda(\lambda_{0})}\approx\langle{\lambda}|$ $e^{\beta_{0}\widetilde{H}_{L,0}}+\left(\beta_{0}\,\delta\widetilde{H}_{L,0}+\delta\beta_{0}\widetilde{H}_{L,0}\right)+\frac{1}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\delta\beta_{0}\,\beta_{0}\widetilde{H}_{L,0}^{2}$ (48)
$+\frac{1}{2}\beta_{0}^{2}\,\delta\widetilde{H}_{L,0}^{2}+\delta\beta_{0}\left(\delta\widetilde{H}_{L,0}+\beta_{0}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\frac{1}{2}\delta\beta_{0}\widetilde{H}_{L,0}^{2}\right)|{\lambda}\rangle\leavevmode\nobreak\ .$
In principle of course, one could choose to keep terms to arbitrarily high order in the perturbation $h$. In this work, we limit ourselves to second order, since this already reveals several corrections relative to previous work.
3.4 Second-order perturbation theory for von Neumann entropy
We now wish to apply our machinery above to compute the change in the von Neumann entropy (27) under a perturbation of the type (38). Again, while we have endeavored to keep the present discussion as general as possible, we will shortly consider the specific case of a double-trace deformation of the TFD in the next section, so the reader may wish to keep that situation in mind if she finds the physical context helpful.
Concretely, denoting the density state in the type II algebras on the right-hand side before and after the deformation, respectively, by $\widehat{\rho}_{R,0}$ and $\widehat{\rho}_{R}$, we wish to compute
$\Delta S=S(\widehat{\rho}_{R})-S(\widehat{\rho}_{R,0})=-\mathrm{tr}\,\widehat{\rho}_{R}\ln\widehat{\rho}_{R}+\mathrm{tr}\,\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}\leavevmode\nobreak\ ,$ (49)
as mentioned at the beginning of subsec. 3.3. Note however that while $\widehat{\rho}_{R}=U\widehat{\rho}_{R,0}U^{\dagger}$, this is not a unitary operation with respect to the right algebra $\widehat{\mathfrak{A}}_{R}$, as explained in subsec. 3.1, so the change in entropy will generically be non-zero. Indeed, from the discussion leading up to (35), this becomes
$\Delta S=-\int\!\mathrm{d}\lambda_{0}\left(e^{\lambda(\lambda_{0})}-e^{\lambda_{0}}\right)\left<{\Psi_{0}}\right|{\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ .$ (50)
It thus remains only to compute the weight $e^{\lambda(\lambda_{0})}$, the meaning of which as an operator expectation value was given to second order in (48). Note that the leading exponential factor will cancel in the final expression for the change in entropy, since
$e^{\lambda_{0}}=\left<{\lambda}\right|{e^{\beta_{0}\widetilde{H}_{L,0}}}\left|{\lambda}\right>\leavevmode\nobreak\ .$ (51)
As we will now see, $\delta\widetilde{H}_{L,0}$ admits an expression as an infinite series of nested commutators. This follows from choosing a form for the unitary $U$ in (38),
$U=e^{-i\tau H}\qquad\mathrm{with}\qquad\tau\coloneqq t-t_{0}\leavevmode\nobreak\ ,$ (52)
where $\tau$ represents the duration in Lorentzian time for which the perturbation is turned on. We then have
$Ue^{\widetilde{H}_{L,0}}U^{\dagger}$ $=\exp\left\{U\widetilde{H}_{L,0}U^{\dagger}\right\}=\exp\left\{e^{-i\tau H}\widetilde{H}_{L,0}\,e^{i\tau H}\right\}$ (53)
$=\exp\left\{\sum_{n=0}^{\infty}\frac{[(-i\tau H)^{n},\widetilde{H}_{L,0}]}{n!}\right\}\leavevmode\nobreak\ ,$
where we have used the Baker-Campbell-Hausdorff (BCH) formula in the form
$e^{A}Be^{-A}=\sum_{n=0}^{\infty}\frac{[(A)^{n},B]}{n!}\leavevmode\nobreak\ ,$ (54)
where $[(A)^{n},B]$ is the iterated commutator
$[(A)^{n},B]\coloneqq\underbrace{[A,\ldots[A,[A}_{\textrm{$n$ times}},B]]\ldots]\leavevmode\nobreak\ ,\qquad\quad[(A)^{0},B]\coloneqq B\leavevmode\nobreak\ .$ (55)
As will become apparent in sec. 4, it will be convenient to break the Hamiltonian $H$ appearing in the unitary deformation into a decoupled part with a formal decomposition in both the left- and right-algebras, $H_{0}=H_{L,0}+H_{R,0}$, and an interacting $\delta H_{0}$ piece that couples the two, so that
$H=H_{0}+\delta H_{0}\leavevmode\nobreak\ ,$ (56)
where by construction, $[H_{0},\widetilde{H}_{L,0}]=0$. Practically, this amounts to replacing $H$ by $\delta H_{0}$ in the BCH expression above:
$Ue^{\widetilde{H}_{L,0}}U^{\dagger}=\exp\left\{\sum_{n=0}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\right\}=\exp\left\{\widetilde{H}_{L,0}+\sum_{n=1}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\right\}\leavevmode\nobreak\ ,$ (57)
and hence
$\delta\widetilde{H}_{L,0}=\sum_{n=1}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\approx-i\tau[\delta H_{0},\widetilde{H}_{L,0}]-\frac{\tau^{2}}{2}[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]\leavevmode\nobreak\ ,$ (58)
where we used the fact that $[\delta H_{0},\widetilde{H}_{L,0}]\sim O(h)$, and in the second step we have expanded to second order in the strength of the perturbation $h$. Thus, inserting this into (48), the perturbative expression for the weight to this order is
[…]
where we have taken $\delta\beta_{0}\approx\beta_{1}+\beta_{2}$ as per (45), and again $\{\mathcal{O}_{1},\mathcal{O}_{2}\}$ denotes the anticommutator. On the first line, everything except the leading exponential term is $O(h)$, while terms on the second and third line are all $O(h^{2})$. Terms carrying factors of $\tau$ depend on the deformation, while those without depend only on the expansion of the inverse temperature (of course, some terms depend on both). Substituting this back into (50), we find that order-by-order, the difference in the von Neumann entropy is
$\Delta S^{(0)}=$ $0\leavevmode\nobreak\ ,$ (60)
$\Delta S^{(1)}=$ $\int\!\mathrm{d}\lambda_{0}\langle{\lambda}|\,i\tau\beta_{0}[\delta H_{0},\widetilde{H}_{L,0}]+\frac{i\tau}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}$
$\qquad\quad\quad-\beta_{1}\widetilde{H}_{L,0}\left(1+\beta_{0}\widetilde{H}_{L,0}\right)|{\lambda}\rangle\left<{\Psi_{0}}\right|{\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ ,$
$\Delta S^{(2)}=$ $\int\!\mathrm{d}\lambda_{0}\langle{\lambda}|\frac{\tau^{2}}{2}\Big{(}\beta_{0}[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]+\frac{1}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]\}$
$\qquad\quad\quad+\beta_{0}^{2}[\delta H_{0},\widetilde{H}_{L,0}]^{2}\Big{)}+i\tau\left(\beta_{1}[\delta H_{0},\widetilde{H}_{L,0}]+\beta_{0}\beta_{1}\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}\right)$
$\qquad\quad\quad-\frac{1}{2}\beta_{1}^{2}\widetilde{H}_{L,0}^{2}-\beta_{2}\widetilde{H}_{L,0}\left(1+\beta_{0}\widetilde{H}_{L,0}\right)|{\lambda}\rangle\left<{\Psi_{0}}\right|{\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ ,$
where $\Delta S^{(n)}$ denotes the change in the generalized entropy at order $n$, i.e., $\Delta S^{(n)}\sim O(h^{n})$. Note that the operators $\mathcal{O}_{L,R}$ will generally not commute with $\widetilde{H}_{L,0}$, so the commutators appearing in these expressions will generally be off-diagonal in the $\lambda$ basis (which is why we could not simply write the weight in terms of eigenvalues as in the naïve formalism on the left-hand side of (30)). Furthermore, the appearance of terms containing $\widetilde{H}_{L,0}$ outside commutation relations requires an explicit choice of Hamiltonian in order to proceed. [^13] As will become more clear in subsec. 4.3 below, the factors of $\widetilde{H}_{L,0}$ within commutators can be trivially transmuted into time derivatives via the Heisenberg equation of motion. Nonetheless, this expression has the advantage of being extremely general, and can in principle be computed in one’s scenario of interest by simply inserting expressions for the first- and second-order commutators appearing here. Additionally, as promised in footnote 11, the state $g(X)$ has again dropped out: the integral depends only on the spectrum of the operators appearing here, not on the state.
In the next section, we apply this general result to the specific case of a double-trace deformation of the TFD, which is dual to a traversable wormhole in the eternal BTZ black hole Gao:2016bin .
4 The traversable wormhole and generalized entropy
In this section, we describe the traversable wormhole of GJW Gao:2016bin in terms of algebras of observables and apply the perturbation theory developed in the previous section to analyze the change in generalized entropy as a consequence of the double-trace deformation
$\delta I=\int\!\mathrm{d}t\,\mathrm{d}^{d}x\,h(t,x)\mathcal{O}_{R}(t,x)\mathcal{O}_{L}(-t,x)\leavevmode\nobreak\ ,$ (61)
(we deviate from GJW in using $I$ for the action of the theory, to avoid confusion with the von Neumann entropy $S$). This deformation renders the wormhole connecting the two boundaries traversable, changing the local structure of observables localized on either side.
As a brief recap before commencing, recall that the thermofield double state of two non-interacting boundary CFTs is holographically dual to the eternal black hole in AdS maldacena:2001kr . In $2\!+\!1$ dimensions, the bulk spacetime is that of the BTZ black hole banados:1992wn (for a review, see carlip:1995qv ), with the metric
$\mathrm{d}s^{2}=-(r^{2}-r_{h}^{2})\,\mathrm{d}t^{2}+(r^{2}-r_{h}^{2})^{-1}\mathrm{d}r^{2}+r^{2}\,\mathrm{d}\phi^{2}\leavevmode\nobreak\ ,$ (62)
where $r_{h}$ is the horizon radius, and we have set $\ell_{P}=1$. By examining the periodicity of the temporal coordinate in Euclidean signature, one finds the inverse temperature of this black hole is $\beta=\frac{2\pi}{r_{h}}$. As an eternal black hole in AdS, it has both a left ($L$) and right ($R$) exterior, excitations about which are described by light primary operators in the respective left or right boundary CFT. It is the algebras of these operators to which we will apply our perturbative machinery below (see Jefferson:2018ksk for an early attempt to study these algebras in this particular context with more details on the bulk/boundary identification, as well as the more recent related works Leutheusser:2021frk ; Leutheusser:2021qhd ).
Turning on the double-trace deformation (61) couples the two boundaries, creating a negative-energy shockwave in the bulk that renders the wormhole traversable for a time controlled by the strength of the deformation $h(t,x)$ Gao:2016bin . An important aspect of this in the present work is that this bilocal interaction introduces non-locality from the perspective of the exterior algebras: elements of the left algebra become elements of the right, and vice versa.
4.1 The semiclassical expansion and area
In order to better place our discussion in the context of previous work, we first briefly review the semiclassical expansion, and the relation between areas and modular charges. The motivation for the former is that we are interested in building up the algebra of observables of propagating gravitons and matter in the bulk. As a side goal, we cement the relation between geometric area and modular Hamiltonians since the area enters the generalized entropy which we ultimately compute.
One way to see the relation between area and modular charges is to realize that the conserved charge associated to boosts about an entangling surface in general relativity is simply the area of that surface Kirklin:2018gcl . A quantum generalization of this statement can be made using the semiclassical expansion of the metric,
$g_{\mu\nu}=g_{\mu\nu}^{(0)}+g_{\mu\nu}^{(1/2)}+g_{\mu\nu}^{(1)}+\,\ldots$ (63)
where the superscript denotes the order of $\hbar$. One then solves the semiclassical Einstein equations,
$G_{\mu\nu}=\left<{T_{\mu\nu}}\right>_{\phi}$ (64)
order-by-order, in a particular state $\phi$. The algebra of observables then contains propagating gravitons, so the modular charge (and hence the corresponding entangling surface) knows about quantum gravitational fluctuations about the classical background. In particular, this means that changes in the state $\phi$ may influence the area. To see this in more detail, we use the relation obtained by Wall in his proof of the generalized second law Wall:2011hj :
$C_{0}-A(\Lambda_{0})=8\pi G\left<K_{0}(\Lambda_{0})\right>_{\phi},$ (65)
where $A(\Lambda_{0})$ is the codimension-2 area of a causal horizon along a spacelike cut $\Lambda_{0}$, and $K_{0}$ is the corresponding boost generator about the bifurcation point, i.e., the modular charge. The constant $C_{0}$ is the area of the cut at infinity $A(\infty)$, which is sensitive to the asymptotics of the state. [^14] Technically, $C_{0}$ governs the asymptotic behavior of an ensemble of states since there are many states with different bulk configurations leading to the same area at infinity. In particular, the modular charge expectation value is order $\hbar$ since it is the term corresponding to quantized fluctuations of the matter fields. This means that the left-hand side should also be $\mathcal{O}(\hbar)$. Thus, the area on the left-hand side is computed using the metric $g_{\mu\nu}$ in the particular state $\phi$, up to $\mathcal{O}(\hbar)$, which implies that the area itself depends on the state explicitly.
Now, imagine applying this relation to the state $U|{\phi}\rangle$, where $U$ is some perturbing interaction. Then
[…]
where in the second equality, we used the fact that the modular charges are related as ${K=UK_{0}U^{\dagger}}$, cf. (33). The cut $\Lambda$ denotes the region of interest in the state $U\phi$ (note that the action of $U$ may not have a geometric interpretation in general). The cancellation of the $U$-dependence between operators and states allows us to to take the difference between (65) and (66) to obtain
$\Delta C-\Delta A=8\pi G\left<{K_{0}(\Lambda)-K_{0}(\Lambda_{0})}\right>_{\phi}.$ (67)
This expression allows one to compute the change in area under unitary transformations in terms of asymptotic data and the original modular charge for the original state. In general, since the QES jumps after the perturbation, the regions defined by $\Lambda$ and $\Lambda_{0}$ are not the same. Furthermore, for the traversable wormhole, the asymptotics before and after the perturbation are not the same as the ADM mass changes as a consequence of the perturbation, and therefore $\Delta C\neq 0$. As we will discuss in the more detailed comparison below, GJW argue that to linear order in the strength of the deformation, the area does not change. However, we can already see from the above expression that at least the constant $C$ must change to linear order.
As an aside, we note that – at least in some cases where the deformation $U$ has a geometric interpretation – one might expect a covariance property of the algebra of observables, and in particular $K_{0}(\Lambda)=K(\Lambda_{0})$ (i.e., transforming a localized operator about $\Lambda$ is the same as undoing the transformation of the operator while shifting the region). In this case, (67) takes the pleasing form
$\Delta A=\Delta C-8\pi G\left<{\delta K_{0}(\Lambda_{0})}\right>_{\phi}\leavevmode\nobreak\ ,$ (68)
where $\delta K_{0}(\Lambda_{0})\coloneqq K(\Lambda_{0})-K_{0}(\Lambda_{0})$. This expression has the nice property that the change does not depend on the explicit shift in the QES: since the difference $\delta K_{0}$ is computed relative to the original cut $\Lambda_{0}$, it is not necessary to know how the QES changes (whereas any dependence on $\Lambda$ requires knowledge of the new bifurcation point).
The relation (67) describes the change in only the leading area term of the generalized entropy, in contrast to our more complete analysis below, but illustrates how areas transform under a deformation using Wall’s relation (65) to the modular charge. Having reviewed the semiclassical expansion governing the quantum gravitational effects included in the algebra of observables, we next turn to the example case of the traversable wormhole, and compute the changes to the generalized entropy under the double-trace deformation more explicitly.
4.2 Deformed algebra of observables
Since the TFD is symmetric, we may work from the perspective of either side; here, we choose the right boundary for concreteness. In keeping with our general notation above, the type III1 algebra of observables $\mathfrak{A}_{R,0}$ before the deformation is spanned by single-trace operators acting on the thermofield double state $|{\Psi_{0}}\rangle$. To construct the crossed product algebra $\widehat{\mathfrak{A}}_{R,0}$, one seeks to adjoin the right modular charge $H_{R,0}$. In large-$N$ gauge theories however, there is a subtlety owing to the fact that the Hamiltonian typically carries an explicit factor of $N$, and is thus not a valid observable Witten:2021unn ; Chandrasekaran:2022cip . To resolve this, one renormalizes and instead adjoins $\widetilde{H}_{R,0}=H_{R,0}/N$, and we denote these suitably-normalized charges by tildes throughout the text. [^15] Note we avoid subtracting the expectation value $H_{R,0}\right>/N$ because this removes the vacuum contribution of interest, as explained in AliAhmad:2023etg . Meanwhile, the dual bulk algebra is comprised of field degrees of freedom localized in the right exterior of the eternal black hole. Adjoining the suitably-renormalized modular charge on the boundary corresponds to adjoining the ADM mass in the bulk, and respects the time-translation isometry of the background spacetime. The $N$-dependence on the boundary corresponds to $G_{\rm N}$-dependence in the bulk, the physical interpretation of which is that the dual gravity picture incorporates geometric fluctuations about the fixed classical background, which one can think of as being generated by propagating gravitons, cf. the semiclassical picture reviewed above. Of course, the story for the left boundary/exterior is precisely identical, and the algebras $\mathfrak{A}_{R,0}$ and $\mathfrak{A}_{L,0}$ are mutual commutants. Formally, the Hamiltonian of the two CFTs on the boundary before the deformation may be written in terms of the (one-sided) modular charges as
$H_{0}=H_{L,0}+H_{R,0}\leavevmode\nobreak\ .$ (69)
Similarly, the corresponding modular Hamiltonian is formally
$H_{\mathrm{mod},0}=H_{R,0}-H_{L,0}\leavevmode\nobreak\ ,$ (70)
and does not contain any mixing between the two sides.
The theory on the boundary after the perturbation will differ from that of the unperturbed theory, and hence the resulting algebra of the perturbed TFD also differs, but – importantly – can be constructed from the original left and right algebras of exterior operators by virtue of the form of the interaction (61). Formally, the boundary Hamiltonian will now contain a bilocal coupling term $\delta H_{0}$,
$H=H_{L}+H_{R}=H_{L,0}+H_{R,0}+\delta H_{0}\leavevmode\nobreak\ .$ (71)
In other words, the perturbed algebras will still satisfy $\mathfrak{A}_{L}^{\prime}=\mathfrak{A}_{R}$, except that now each single-sided algebra will encode information from both $\mathfrak{A}_{L,0}$ and $\mathfrak{A}_{R,0}$. The corresponding bulk picture is illustrated in fig. 1. As we will show below, one can formally split the coupling $\delta H_{0}$ into corrections to the left and right charges, $\delta H_{L,0}+\delta H_{R,0}=\delta H_{0}$, so that, e.g., $H_{R}=H_{R,0}+\delta H_{R,0}$ (and similarly for $H_{L}$), but these are not localized to their respective algebras. Rather, $\delta H_{R,0}$ depends on operators $\mathcal{O}_{L}\in\mathfrak{A}_{L,0}$, and conversely for $\delta H_{L,0}$. [^16] We suppress the subscript 0 on operators in the unperturbed algebras, since all operators appearing in this work will always belong to $\mathfrak{A}_{0}$ rather than $\mathfrak{A}$. In this case, one expects the modular Hamiltonian to also pick up a bilocal interaction term, so that
$H_{\mathrm{mod}}=H_{\mathrm{mod},0}+\delta H_{\mathrm{mod},0}\leavevmode\nobreak\ .$ (72)
Non-perturbatively, this is precisely (21), and is indeed the modular Hamiltonian of the algebra $\mathfrak{A}$ associated to the perturbed TFD state.
Figure 1: (Left) Unperturbed TFD state, with the spacetime regions to which the bulk exterior algebras are associated shaded in blue. These are holographically dual to the boundary algebras $\mathfrak{A}_{L,0}$, $\mathfrak{A}_{R,0}$. (Right) TFD after the double-trace deformation, which injects negative energy into the black hole, thereby decreasing the area; the new horizons are shown in red, along with the corresponding exterior regions dual to the boundary algebras $\mathfrak{A}_{L}$, $\mathfrak{A}_{R}$. Importantly, each of these will contain degrees of freedom from both $\mathfrak{A}_{L,0}$ and $\mathfrak{A}_{R,0}$ due to the traversability of the wormhole, i.e., the causal connectivity between the original and deformed exteriors.
Slightly more concretely, suppose we turn on the double-trace deformation (61) at a time $t_{0}$, and evolve the TFD state with the corresponding interacting Hamiltonian (71) until time $t>t_{0}$; the state at time $t$ is then
$|{\Psi}\rangle\coloneqq U(t_{0},t)|{\Psi_{0}}\rangle=e^{-i(t-t_{0})H}|{\Psi_{0}}\rangle\leavevmode\nobreak\ ,$ (73)
where $U(t_{0},t)\coloneqq e^{-i(t-t_{0})H}$, cf. (52). The type III1 algebra on the right boundary $\mathfrak{A}_{R}$ is still built out of single-trace operators acting on $|{\Psi}\rangle$; and to obtain the corresponding crossed product algebra $\widehat{\mathfrak{A}}_{R}$, one needs to adjoin the right modular charge $H_{R}$. However, as just discussed, this is no longer $H_{R,0}$: rather, the new modular charge is $H_{R}=UH_{R,0}U^{\dagger}$, and contains operators that were initially localized in the left algebra, $\mathcal{O}_{L}\in\mathfrak{A}_{L,0}$ due to the form of the interaction (61). Abstractly however, the normalization issue in the large-$N$ limit is the same as before, and upon completing the crossed product construction on obtains the type II algebra $\widehat{\mathfrak{A}}_{R}$ of the right boundary; similarly for $\widehat{\mathfrak{A}}_{L}$.
Of course, the decomposition (70) is purely formal because there is no factorization between $R$ and $L$ in the type III theory, i.e., the one-sided charges do not generate well-defined states in their respective algebras. Nonetheless, in the unperturbed theory, these objects are integrals over local operators on two spacelike separated regions, and hence we may still say that $[H_{R,0},H_{L,0}]=0$. Moreover, in the crossed product algebra, the modular operator admits a genuine factorization between $R$ and $L$. Thus, it is well-motivated to consider
$=[UH_{R,0}U^{\dagger},UH_{L,0}U^{\dagger}]=U[H_{R,0},H_{L,0}]U^{\dagger}=0\leavevmode\nobreak\ ,$ (74)
even though $H_{R/L}$ has non-trivial support on $L/R$. This will translate to a decomposition of the modular operator in the corresponding crossed product algebra of observables.
Incidentally, note that the relevant question that determines the extent of the new exterior regions is where the quantum extremal surface (QES) shifts as a result of the deformation. That is, in the unperturbed state, the QES sits at the bifurcation surface (the intersection of the blue horizons in the left panel of fig. 1). In the previous paper Jefferson:2018ksk applying modular theory to this scenario, the deformed exteriors were modeled as overlapping, effectively implying two solutions for the QES, located at points $1_{\mathrm{l}}$ and $1_{\mathrm{r}}$ in fig. 2, which bound the right and left exterior regions, respectively. While this is consistent with the symmetry of the system from the perspective of complementarity (in the spirit of Susskind:1993if ), it appears in tension with the purity of the state, which suggests that the QES instead moves directly upwards, to the point labeled 2 in fig. 2, as was also argued in GJW Gao:2016bin . To the best of our knowledge, this has not been explicitly proven, but could in principle be done for the BTZ black hole by matching geometries as in Hirano:2019ugo and extremizing the area functional.
Figure 2: Close-up of the center region in the right panel of fig. 1, showing various intersections of horizons. The QES begins at the original bifurcation surface labeled 0 in the unperturbed TFD. After the deformation, the QES may either move directly upwards to the point labeled 2, or split into two solutions at the points labeled $1_{\mathrm{l,r}}$. Note that in the latter case, the left and right boundary algebras overlap in the central diamond region. From the perspective of black hole complementarity Susskind:1993if , both of these scenarios are consistent with the symmetry of the TFD. However, while the latter has an interesting geometric interpretation in terms of modular inclusions explored in Jefferson:2018ksk , here we follow GJW Gao:2016bin in assuming that the QES moves directly upwards to point 2, which is more obviously consistent with the purity of the TFD. We note however that this does not preclude corrections to the area of the horizon, as we will see below.
4.3 Deformed generalized entropy
In Chandrasekaran:2022cip , it was shown that the von Neumann entropy of the type II algebra $\widehat{\mathfrak{A}}$ in a classical-quantum state is simply the generalized entropy of the region (2),
$S_{\rm gen}(\widehat{\rho})=\frac{A}{4G_{\rm N}}+S_{\rm matter}$ (75)
where $A$ is the area of the underlying region’s boundary, given by the QES, and $S_{\rm matter}$ is the von Neumann entropy of matter fields in the region. In a large-$N$ gauge theory, the first term is $O(N^{2})$, while the second term is $O(1)$. In the semiclassical expansion above, there are also quantum gravitational effects that are suppressed by additional powers of $1/N$. As we will see, our perturbative approach reveals corrections to both the area and the matter terms in this expression. It is important to note however that the fundamental quantity is the von Neumann entropy, and the factorization into distinct parts as in (75) must be taken with a grain of salt non-perturbatively AliAhmad:2024saq . Thus, our main result is the effect of the perturbation on the generalized entropy as a whole, which one can subsequently interpret as changes to the area vs. matter contributions.
Given the setup in the preceding subsections, the only objects required to apply our general analysis to obtain the change in the generalized entropy (75) under the deformation are the first- and second-order commutators $[\delta H_{0},\widetilde{H}_{L,0}]$ and $[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]$. Since ${\delta H_{0}=-\delta S}$ with (61), we have simply
$=-\!\int\!\mathrm{d}t_{1}\,\mathrm{d}^{d}x\,h(t_{1},x)\mathcal{O}_{R}(t_{1},x)[\mathcal{O}_{L}(-t_{1},x),\widetilde{H}_{L,0}]$ (76)
$=-\frac{i}{N}\!\int\!\mathrm{d}t_{1}\,\mathrm{d}^{d}x\,h(t_{1},x)\mathcal{O}_{R}(t_{1},x)\,\partial_{t_{1}}\mathcal{O}_{L}(-t_{1},x)\leavevmode\nobreak\ .$
where we used the fact that in the Heisenberg picture, $[\widetilde{H},\mathcal{O}]=-\frac{i}{N}\partial_{t}\mathcal{O}$. Note that this is first-order in the strength of the deformation due to the explicit factor of $h$, and provides the linear contribution to $\delta H_{L,0}$ in the sense that
$H_{L}=UH_{L,0}U^{\dagger}\approx H_{L,0}-i(t-t_{0})[\delta H_{0},H_{L,0}]\leavevmode\nobreak\ .$ (77)
Similarly, the second-order commutator is
$[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]=-\frac{i}{N}\!\int\!\mathrm{d}t_{1}\mathrm{d}t_{2}\mathrm{d}^{d}x\mathrm{d}^{d}y\,h(t_{1},x)h(t_{2},y)[\mathcal{O}_{R}(t_{2},y)\mathcal{O}_{L}(-t_{2},y),\,\mathcal{O}_{R}(t_{1},x)\dot{\mathcal{O}}_{L}(-t_{1},x)]\leavevmode\nobreak\ ,$ (78)
where the dot denotes the time derivative. In principle, one could compute these using Green functions in the BTZ background, cf. Hirano:2019ugo , and the first-order commutator (76) was also given in Gao:2016bin . However, such a detailed computation is not the goal of the present work; instead, we wish to compare and contrast the perturbation theory analysis in the crossed product with the original work of GJW Gao:2016bin .
Under the deformation, the algebra $\mathfrak{A}_{R}$ localized in the new right exterior contains part of the old interior. Since the perturbation changes the location of the QES, we expect both terms in the generalized entropy (75) to be affected: (1) the area should be the area of the new bifurcation surface (point 2 in fig. 2), and (2) the matter entropy now gains contributions from correlations between exterior and interior operators from the perspective of undeformed theory. In Gao:2016bin , GJW argued heuristically that the area of the QES should not change at leading order in $h$ because “the geometry near the bifurcation is unaffected by the perturbation.” This would amount to setting $\beta_{1}=0$ in (45). A priori however, we see no reason why the area cannot change at leading order; on the contrary, for the wormhole to become traversable, we know that the QES must shift, so the statement that the geometry near the bifurcation surface is unaltered appears puzzling. For this reason, we have kept $\beta_{1}\neq 0$ in our general analysis above, and we will discuss the resulting differences relative to GJW presently. In their work, it was instead argued that the only change to linear order should be in the subleading matter term, and that by the first law of entanglement entropy Blanco:2013joa , this is given by $\delta S_{\mathrm{bulk}}=\beta\delta H_{R,0}$ (in our notation).
Let us compare this to our results. At first order in $h$, we found
$\Delta S_{\mathrm{gen}}^{(1)}=$ $\int\!\mathrm{d}\lambda_{0}\langle{\lambda}|\,\underbrace{i\tau\beta_{0}[\delta H_{0},\widetilde{H}_{L,0}]-\beta_{1}\widetilde{H}_{L,0}}_{O(1/N)}$ (79)
$\qquad\quad\quad+\underbrace{\frac{i\tau}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}-\beta_{0}\beta_{1}\widetilde{H}_{L,0}^{2}}_{O(1/N^{2})}|{\lambda}\rangle\left<{\Psi_{0}}\right|{\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ ,$
(this is just (60), reproduced for convenience, except that we have grouped terms order-by-order in $1/N$, and added the subscript gen in light of the identification of the type II von Neumann entropy with generalized entropy mentioned above), which one can in principle evaluate by substituting in (76) and (78). The first term, with the commutator $[\delta H_{0},\widetilde{H}_{L,0}]$, corresponds to the contribution that GJW Gao:2016bin identified from the first law of entanglement entropy, cf. (77). The others are new, and while they are all $O(h)$, one appears at $O(1/N)$, while the other two are $O(1/N^{2})$ due to the quadratic appearance of the normalized left modular charge. These effects vanish in the $N\to\infty$ limit, and hence should be understood as quantum corrections. Notably however, while the $O(1/N^{2})$ term proportional to $\tau$ may be interpreted as a correction to $S_{\mathrm{matter}}$, both terms proportional to $\beta_{1}$ manifestly contribute to the area term $A/4G$, since they explicitly depend on the linear correction to the horizon temperature. In other words, the von Neumann entropy computed via the crossed product knows about both the area and matter terms in the generalized entropy, and naturally includes corrections to both.
In addition, we have also computed the contribution at order $h^{2}$:
$\Delta S_{\mathrm{gen}}^{(2)}=$ $\int\!\mathrm{d}\lambda_{0}\langle{\lambda}|\underbrace{\frac{\tau^{2}}{2}\beta_{0}[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]+i\tau\beta_{1}[\delta H_{0},\widetilde{H}_{L,0}]-\beta_{2}\widetilde{H}_{L,0}}_{O(1/N)}$ (80)
$\qquad\quad\quad+\frac{\tau^{2}}{4}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]\}+\frac{\tau^{2}}{2}\beta_{0}^{2}[\delta H_{0},\widetilde{H}_{L,0}]^{2}$
$\qquad\quad\quad\underbrace{+\leavevmode\nobreak\ i\tau\beta_{0}\beta_{1}\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}-\frac{1}{2}\beta_{1}^{2}\widetilde{H}_{L,0}^{2}-\beta_{0}\beta_{2}\widetilde{H}_{L,0}^{2}\leavevmode\nobreak\ }_{O(1/N^{2})}|{\lambda}\rangle\left<{\Psi_{0}}\right|{\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ ,$
which one can again evaluate by substituting in the first- and second-order commutators above, though as mentioned before, one would additionally require a particular choice for the Hamiltonian on the boundary due to free factors of $\widetilde{H}_{L,0}$ that cannot obviously be transmuted into time derivatives. Observe that this expression contains three terms at $O(1/N)$, as well as five terms at $O(1/N^{2})$. Furthermore, the last two terms on the first line (at order $1/N$), as well as all three terms on the last line (at order $1/N^{2}$) explicitly involve subleading or sub-subleading terms in the expansion of the inverse temperature (45), and thus represent quantum corrections to the horizon area when splitting the von Neumann entropy in the form (75).
To summarize: by first transmuting the type III algebra of exterior operators to a type II algebra by adjoining the modular charge, we have perturbatively computed the change in the generalized entropy under a double trace deformation of the TFD to second order in both $h$ and $1/N$. The change in $\beta$ arises from the shift in the horizon radius, which changes the area term in the generalized entropy (75), and the $O(1/N^{2})$ terms represent sub-subleading contributions from quantum gravitational effects in the bulk. In addition to the greater rigour and control relative to previous heuristic arguments, we emphasize that our approach allows one to compute in principle arbitrarily high orders in perturbation theory, since the weight $e^{\lambda(\lambda_{0})}$ ultimately amounts to a straightforward Baker-Campbell-Hausdorff expansion.
5 Discussion: the fate of the algebraic approach in quantum gravity
While a classic tool in the theory of operator algebras, the crossed product has only recently emerged in high-energy theory as a tool for accessing semifinite features of algebras of observables in quantum field theory and gravity Witten:2021unn ; Chandrasekaran:2022cip ; Chandrasekaran:2022eqq ; Jensen:2023yxy ; Klinger:2023tgi ; AliAhmad:2023etg ; AliAhmad:2024eun , as well as a means towards a fundamental description of holographic theories away from the semiclassical limit AliAhmad:2024wja ; AliAhmad:2024vdw ; DeVuyst:2024uvd ; AliAhmad:2024saq . In light of this, it is only natural to wonder if the crossed product is stable under perturbations, that is: are the crossed product algebras before and after a transformation related in a tractable way?
In this work, we give an affirmative answer to this question in the case of unitary [^17] Technically, our results would also apply to the more general case when $U$ is merely orthogonal, but the physical motivation for this is less clear. transformations of the full algebra. We construct a perturbation theory built around the crossed product that allows us to describe changes in the observable, state, and expectation value structure of a theory under a unitary transformation. Just as there exists an open quantum system description of a total unitary that induces an operation on its subsystems Lidar:2019qog , the crossed product can model the changes in observables as a consequence of a total unitary on the algebra and its commutant that does not behave like a unitary on either one. [^18] For the relation between open quantum thermodynamics and the crossed product, we refer the reader to Cirafici:2024ccs ; Cirafici:2024jdw . We give explicit general formulas for the changes in entropy under a unitary transformation, and specialize to the case of a bilocal deformation as a situation of physical interest.
The initial impetus for this investigation was the intrigue of the black hole interior Papadodimas:2012aq ; Papadodimas:2013jku ; Verlinde:2012cy ; Verlinde:2013qya ; Verlinde:2013uja ; Papadodimas:2017qit ; Maldacena:2013xja ; Jefferson:2018ksk , as well as black hole evaporation. In the eternal black hole, the algebras localize on the causal development of either boundary, and the interior is causally disconnected. After turning on the double-trace deformation that renders the wormhole traversable, the causal development of one boundary seems to include part of the original interior. Concretely, one sees this as a mixture of the original left and right degrees of freedom on the new one-sided algebras. For a single-sided evaporating black hole, one might consider the left and right as the black hole interior and its radiation, respectively, insofar as under the evaporation process modeled as a unitary on the total system, the radiation degrees of freedom will inexorably mix with those of the black hole. This points to a degree of non-locality in the evaporation process that resembles some of the toy models in the current literature proposing that the black hole interior is part of the late time radiation (see Penington:2019npb ; Almheiri:2019psf ; Almheiri:2020cfm and subsequent works). The difference is that in those models, this identification arises via an ad hoc choice of cut when performing the path integral (i.e., one declares a density matrix with support in two disconnected regions), whereas here it arises naturally as a consequence of the geometric interpretation of the deformation that, at least infinitesimally, may be considered as an analogue of the evaporation process.
Expanding our perspective further, our work reveals a fundamental issue in the algebraic approach to quantum field theory and gravity: how does one talk about algebras of observables on regions that are not fixed? In quantum gravity, it is expected for regions not only to fluctuate but also transform into others and change their geometry. The former is a perturbative statement while the other is non-perturbative. The algebraic formalism used here presupposes a fixed region (albeit one which undergoes a controlled deformation) in which the field degrees of freedom localize. Can the algebraic approach be made to accommodate geometric and topological changes in spacetime subregions? In the context of holography, the answer tentatively seems to be yes, insofar as the boundary provides a fundamental description for the theory where subregions may be defined for semiclassical states and their fluctuations captured by excitations. Upon the addition of heavier boundary operators however, the semiclassical description may break down and the bulk may no longer admit a geometric description. This problem is also circumvented in lower dimensions, where the fundamental description of JT gravity is given by a matrix model that makes no reference to geometry Saad:2019lba . In the bulk, the picture is less clear, and we hope that our perturbative analysis provides a step towards the historied challenge of combining algebraic quantum field theory and quantum gravity Stewart:1974uz ; Fredenhagen:1989kr ; Haag:1990ht ; Brunetti:2013maa ; Rejzner:2016yuy ; Brunetti:2016hgw ; Brunetti:2022itx .
Acknowledgements
S.A.A. would like to thank Ahmed Almheiri and Simon Lin for helpful discussions.